# ConSmax: Hardware-Friendly Alternative Softmax with Learnable Parameters

### 31 Jan 2024 v1

@@ Abstract
The self-attention mechanism sets transformer-based large language model (LLM) apart from the convolutional and recurrent neural networks. Despite the performance improvement, achieving real-time LLM inference on silicon is challenging due to the extensively used Softmax in self-attention. Apart from the non-linearity, the low arithmetic intensity greatly reduces the processing parallelism, which becomes the bottleneck especially when dealing with a longer context. To address this challenge, we propose Constant Softmax (ConSmax), a software-hardware co-design as an efficient Softmax alternative. ConSmax employs differentiable normalization parameters to remove the maximum searching and denominator summation in Softmax. It allows for massive parallelization while performing the critical tasks of Softmax. In addition, a scalable ConSmax hardware utilizing a bitwidth-split look-up table (LUT) can produce lossless non-linear operation and support mix-precision computing. It further facilitates efficient LLM inference. Experimental results show that ConSmax achieves a minuscule power consumption of 0.43 mW and area of 0.001 mm at 1-GHz working frequency and 22-nm CMOS technology. Compared to state-of-the-art Softmax hardware, ConSmax results in 14.5 energy and 14.0 area savings with a comparable accuracy on a GPT-2 model and the WikiText103 dataset.
@@
Transformer-based LLM models [1, 2] have served as the foundation across a broad range of machine learning fields, from natural language processing [3, 4, 5] to computer vision [6, 7], etc. The enviable improvement can be attributed to the unique self-attention mechanism. Different from previous convolutional or recurrent network (CNN and RNN), this mechanism enhances LLMs to capture information across input contexts (i.e. tokens) regardless of their distance. The fabulous prospects in LLM have also driven efforts on specific accelerator design [8, 9, 10, 11]. However, accelerating LLM inference on silicon is challenging due to its low arithmetic intensity. It prevents the further applications of LLM, especially on edge devices.
Softmax is the main factor exacerbating the inefficiency of LLM. Figure 1 details the LLM structure as well as the self-attention structure. The self-attention first transform the input token embeddings into separate representations termed Query (Q), Key (K) and Value (V). The inner product is then computed between Q and K to determine the attention scores (S), reflecting the relevance between each pair of tokens. These scores are normalized as probabilities (P) and used to calculate weighted averages of V. The averaged result comprises the self-attention output. In addition to the mentioned matrix multiplications, the Softmax is the pivot to convert attention scores to probabilities. To compute the Softmax, the entire score vector should be allocated and iterated over to calculate the maximum score and score summation. Therefore, the after-Softmax PV should be blocked until the before-Softmax QK and Softmax finish. Due to the low parallelism opportunity, Softmax reduces the hardware utilization of main-steam GPUs and TPUs, as well as the potential hardwares. Softermax [12] shows that the Softmax can lead more than 30% overhead of LLM latency, especially at the longer input sequence found in state-of-the-art LLM models [2, 13, 14].
@@
In summary, previous Softmax-oriented works suffer from inefficiency for lack of high parallelism and accuracy. To address these challenges, there is a noteworthy observation that the the maximum score can be replaced by arbitrary values to scale the numerator and denominator in Softmax. In addition, the probability vector does not have to be a unit vector to ensure LLM accuracy. Therefore, the score summation to generate the Softmax denominator can be avoided . Inspired by the above insights, this paper proposes Constant Softmax (ConSmax), a software-hardware co-design as an efficient Softmax alternative. Marking a departure from prior approaches, ConSmax can increase computing parallelism and produce lossless normalization for LLM models. In short, the key contributions are listed as follows.
@@
We extensively evaluate ConSmax on a GPT-2 model and the WikiText103 [39] dataset. Experimental results show that ConSmax achieves a minuscule power consumption of 0.43 mW and area of 0.001 mm at 200-MHz working frequency and 22-nm CMOS technology. Compared to state-of-the-art Softmax hardware, ConSmax results in more than 15 energy and 10 area savings.
@@ III ConSmax Algorithm
Although Softmax operation accounts for a relatively small portion of the overall self-attention layer, Softmax is much more challenging to design into efficient hardware than matrix-multiplication. Therefore, it can incur significant overhead if not handled appropriately. In section, we propose ConSmax algorithm, which can avoid data synchronization and hence increase computing parallelism in Softmax.
@@ III-A Convert Softmax to ConSmax
The Softmax operation in Figure 3(a) requires the entire score vector to calculate the maximum score and score summation. However, it is unreasonable to generate all scores at the same cycle, especially since the current LLM models allow thousands to hundreds of thousands of input tokens [2, 13, 14] (such as 8K to 128K tokens). Therefore, the potential hardware should buffer the exponent results for all score elements before Softmax can proceed. Note that the exponent operation would explode the numerical range. It results in more memory consumption to maintain the precision of these intermediate results. To increase Softmax parallelism, the key problem is how to compute each partial Softmax result without requiring maximum score and summation results from other partial Softmax computations.
@@ IV ConSmax Hardware
Based on the algorithm in Section III, we delve into the intricacies of the customized ConSmax hardware. The bitwidth-split LUT architecture can produce lossless non-linear functions and scale for mix-precision computing. In addition, we also detail how to integrate the ConSmax hardware into state-of-the-art transformer accelerators. Thanks to the synchronization-free property, ConSmax can increase parallelism in both LLM summarization and generation stages.
@@ IV-B Integrate ConSmax Hardware to Transformer Accelerator
ConSmax | Softmax | |||||
Process | 22nm | 130nm | 22nm | 130nm | ||
EDA Tool | Design Compiler | OpenRoad | Design Compiler | OpenRoad | ||
Frequency | 1GHz | 200MHz | 1GHz | 200MHz | ||
Power (mW) | 0.425 | 1.82 | 2.69 | 6.18 | 42.2 | 51 |
Area (mm) | 0.001 | 0.007 | 0.015 | 0.014 | 0.18 | 0.2 |
Energy per op(J) | 0.425p | 9.10p | 13.45p | 6.18p | 211.00p | 255.00p |
@@ V-A Experiment Setup
For software performance, we evaluate ConSmax on a small GPT-2 model [41, 42] and the WikiText103 [39] dataset. The benchmark model contains 6 transformer layers with 6 self-attention heads, where the size of embedding features is set to 384. The default context length is set to 256. We substitute RMSNorm [43] for LayerNorm. We train for 30,000 iteration steps. For ConSmax evaluation, we replace Softmax in each of the attention blocks with ConSmax. We report validation loss for the language generation task. This metric measures the ability of the transformer language model on text generation. A lower validation loss value indicates better performance.
A ConSmax prototype is designed under 22nm and 130nm CMOS technology with Verilog RTL. We synthesize ConSmax with the Synopsys Design Compiler and open-source OpenRoad EDA toolchains. The generated netlist is used to evaluate the power and area consumption.
@@ V-B Software Performance
In Figure 6, we conducted a comparative analysis of the variation losses associated with Softmax and the proposed ConSmax. Initially, it is observed that the variation loss for the ConSmax is marginally higher in comparison to Softmax. Nevertheless, this discrepancy diminished over time. After approximately 2000 iterations, the variation losses for Softmax and ConSmax models converge, demonstrating a similar trend in their performance metrics. There is only a 2.3% loss degeneration for ConSmax after the first 2000 iterations and less than 0.9% after the first 10000 iterations.
@@ V-C Hardware Performance
We have carried out a comparative study of our proposed ConSmax against the baseline Softmax. The power and area consumption are evaluated under with 22nm and 130nm CMOS. In order to render a thorough comparison, we have established the testing environment in both commercial Synopsys Design Compiler and open-source OpenRoad EDA tools. Table I summarizes ConSmax hardware performance and compares it with Softmax hardware. Under 22nm CMOS technology, 1-GHz working frequency and 0.9-V power supply, one ConSmax hardware only consumes 0.425mW power and 0.001mm area. Compared to the Softmax counterpart, ConSmax achieves 14.5 power and 14.0 area savings. With a lower 200-MHz working frequency and 130nm CMOS, ConSmax achieves 23.2 power and 25.7 area savings. Note that the OpenRoad tools don’t support synthesis with 22nm CMOS technology, and so we report the performance under 130nm CMOS. With OpenRoad, the ConSmax still shows similar performance improvement by 18.9 power and 13.3 area reductions.
@@ VI Conclusion
This paper presents ConSmax, a software-hardware co-design for efficient Softmax acceleration. ConSmax improves Softermax computing parallelism by avoiding the data synchronization, which produces maximum score and score summation. In addition, the bitwidth-split ConSmax hardware is lossless and scalable to calculating non-linear functions. The experiments show that ConSmax achieves a minuscule power consumption of 0.43 mW and area of 0.001 mm at 1-GHz working frequency and 22-nm CMOS technology. Compared to state-of-the-art Softmax hardware, ConSmax results in 14.5 energy and 10 area savings with a comparable accuracy on GPT-2 model and the WikiText103 dataset.

### 20 Feb 2024 v2

@@ Abstract
The self-attention mechanism sets transformer-based large language model (LLM) apart from the convolutional and recurrent neural networks. Despite the performance improvement, achieving real-time LLM inference on silicon is challenging due to the extensively used Softmax in self-attention. Apart from the non-linearity, the low arithmetic intensity greatly reduces the processing parallelism, which becomes the bottleneck especially when dealing with a longer context. To address this challenge, we propose Constant Softmax (ConSmax), a software-hardware co-design as an efficient Softmax alternative. ConSmax employs differentiable normalization parameters to remove the maximum searching and denominator summation in Softmax. It allows for massive parallelization while performing the critical tasks of Softmax. In addition, a scalable ConSmax hardware utilizing a bitwidth-split look-up table (LUT) can produce lossless non-linear operation and support mix-precision computing. It further facilitates efficient LLM inference. Experimental results show that ConSmax achieves a minuscule power consumption of 0.2 mW and area of 0.0008 mm at 1250-MHz working frequency and 16-nm CMOS technology. Compared to state-of-the-art Softmax hardware, ConSmax results in 3.35 power and 2.75 area savings with a comparable accuracy on a GPT-2 model and the WikiText103 dataset.
@@
Transformer-based LLM models [1, 2] have served as the foundation across a broad range of machine learning fields, from natural language processing [3, 4, 5] to computer vision [6, 7], etc. The enviable improvement can be attributed to the unique self-attention mechanism. Different from previous convolutional or recurrent network (CNN and RNN), this mechanism enhances LLMs to capture information across input contexts (i.e. tokens) regardless of their distance. The fabulous prospects in LLM have also driven efforts on specific accelerator design [8, 9, 10, 11]. However, accelerating LLM inference on silicon is challenging due to its low arithmetic intensity. It prevents the further applications of LLMs, especially on edge devices.
Softmax is the main factor exacerbating the inefficiency of LLMs for long contexts. Figure 1 details the LLM structure as well as the self-attention structure. The self-attention first transform the input token embeddings into separate representations termed Query (Q), Key (K) and Value (V). The inner product is then computed between Q and K to determine the attention scores (S), reflecting the relevance between each pair of tokens. These scores are normalized as probabilities (P) and used to calculate weighted averages of V. The averaged result comprises the self-attention output. In addition to the mentioned matrix multiplications, the Softmax is the pivot to convert attention scores to probabilities. To compute the Softmax, the entire score vector should be allocated and iterated over to calculate the maximum score and score summation. Therefore, the after-Softmax PV should be blocked until the before-Softmax QK and Softmax finish. Due to the low parallelism opportunity, Softmax reduces the hardware utilization of main-steam GPUs and TPUs, as well as the potential hardwares. Softermax [12] shows that the Softmax can lead more than 30% overhead of LLM latency, especially at the longer input sequence found in state-of-the-art LLM models [2, 13, 14].
@@
In summary, previous Softmax-oriented works suffer from inefficiency for lack of high parallelism and accuracy. To address these challenges, there is a noteworthy observation that the the maximum score can be replaced by arbitrary values to scale the numerator and denominator in Softmax. In addition, the probability vector does not have to be a unit vector to ensure LLM accuracy. Therefore, the score summation to generate the Softmax denominator can be avoided . Inspired by the above insights, this paper proposes Constant Softmax (ConSmax), a software-hardware co-design as an efficient Softmax alternative. Marking a departure from prior approaches, ConSmax can increase computing parallelism and produce lossless normalization for LLM models. In short, the key contributions are listed as follows:
@@
We extensively evaluate ConSmax on a GPT-2 model and the WikiText103 [39] dataset. Experimental results show that ConSmax achieves a minuscule power consumption of 0.2 mW and area of 0.0008 mm at 1250-MHz working frequency and 16-nm CMOS technology. Compared to state-of-the-art Softmax hardware, ConSmax results in more than 3.35 power and 2.75 area savings.
@@ III ConSmax Algorithm
Although Softmax operation accounts for a relatively small portion of the overall self-attention layer, Softmax is much more challenging to design into efficient hardware than matrix-multiplication. Therefore, it can incur significant overhead if not handled appropriately. In this section, we propose the ConSmax algorithm, which helps avoid data synchronization and, hence, increase computing parallelism in Softmax.
@@ III-A Convert Softmax to ConSmax
The Softmax operation in Figure 3(a) requires the entire score vector to calculate the maximum score and score summation. However, generating all scores at the same cycle is unreasonable, especially since the current LLM models allow thousands to hundreds of thousands of input tokens [2, 13, 14] (such as 8K to 128K tokens). Therefore, the potential hardware should buffer the exponent results for all score elements before Softmax can proceed. Note that the exponent operation would explode the numerical range. It results in more memory consumption to maintain the precision of these intermediate results. To increase Softmax parallelism, the key problem is how to compute each partial Softmax result without requiring maximum score and summation results from other partial Softmax computations.
@@ III-B Comparison with Partial Softmax
| Proprietary EDA | ConSmax | Softermax | Softmax | Consmax | Softermax | Softmax |
| Process | 16nm | 130nm | ||||
| Max Frequency (MHz) | 1250 | 1111 | 909 | 666.67 | 333.33 | 285.71 |
| Area (mm)a | 0.0008 | 0.0022 | 0.011 | 0.007 | 0.029 | 0.18 |
| Power (mW)b | 0.2 | 0.67 | 1.5 | 2.69 | 8.5 | 51 |
| Optimum Energy per op(pJ) | 0.2 | 0.7 | 1.5 | 4 | 25.5 | 178.5 |
| Opensource EDA | ConSmax | Softermax | Softmax | Consmax | Softermax | Softmax |
| Process | 16nm | 130nm | ||||
| Max Frequency (MHz) | 2000 | 1000 | 500 | 166.67 | 142.86 | 87.72 |
| Area (mm)a | 0.0009 | 0.0019 | 0.011 | 0.015 | 0.033 | 0.2 |
| Power (mW)b | 0.683 | 2.82 | 10.2 | 1.82 | 10.5 | 42.2 |
| Optimum Energy per op(pJ) | 0.3 | 1.4 | 2.7 | 16.7 | 73.5 | 255.3 |
@@ IV ConSmax Hardware
Based on the algorithm in Section III, we delve into the intricacies of the customized ConSmax hardware. The bitwidth-split LUT architecture can produce lossless non-linear functions and scale for mix-precision computing. In addition, we also detail how to integrate the ConSmax hardware into state-of-the-art transformer accelerators. Thanks to the synchronization free property, ConSmax can increase parallelism in both LLM summarization and generation stages.
@@ V-A Experiment Setup
For software performance, we evaluate ConSmax on a small GPT-2 model [41, 42] and the WikiText103 [39] dataset. The benchmark model contains 6 transformer layers with 6 self-attention heads. Embedding features size is set to 384. The default context length is set to 256. We substitute RMSNorm [43] for LayerNorm in post-norm. We train for 30,000 iteration steps. For the ConSmax evaluation, we replaced Softmax in each attention block with ConSmax. We report validation loss for the language generation task. This metric measures the ability of the transformer language model on text generation. A lower validation loss value indicates better performance.
A ConSmax prototype in Verilog is implemented in 16nm and 130nm CMOS technology. We synthesize ConSmax with both proprietary Design Compiler and open-source EDA toolchains. The generated netlist is used to evaluate the power and area consumption.
@@ V-B Software Performance
In Figure 6, we compared the variation losses associated with Softmax and the proposed ConSmax. Initially, it is observed that the variation loss for the ConSmax is marginally higher compared to Softmax. Nevertheless, this discrepancy diminished over time. After approximately 2000 iterations, the variation losses for Softmax and ConSmax models converge, demonstrating a similar trend in their performance metrics. There is only a 2.3% loss degeneration for ConSmax after the first 2000 iterations and less than 0.9% after the first 10000 iterations.
@@ V-C Hardware Performance
We have carried out a comparative study of our proposed ConSmax against the baseline Softmax and Softermax[12]. The power and area consumption are evaluated under 16nm and 130nm CMOS. In order to render a thorough comparison, we have established the testing environment in both commercial EDA and open-source EDA tool.
The feature size for Softmax, Softermax and Consmax are set to 8 for all the experiments.
Table I summarizes ConSmax hardware performance and compares it with Softmax and Softermax hardware. Under 16nm CMOS technology, 500-MHz working frequency and 0.8-V power supply, one ConSmax hardware only consumes 0.2mW power and 0.0008mm area. Compared to the Softermax counterpart, ConSmax achieves 3.35 power, and 2.75 area savings. Compared to Softmax counterpart, Consmax achieves 7.5 power, and 13.75 area savings. With a lower 80-MHz working frequency and 130nm CMOS, ConSmax achieves 3.2 power and 4.14 area savings compared to Softermax. As well as 23.2 power and 25.7 area savings compared to Softmax With opensource EDA, the ConSmax shows similar performance improvement by 4.13 power and 2.11 area reductions.
Figure 8 presents a comparison of the silicon area utilized by three different designs when operated at their maximum frequencies. In comparison with Softermax and Softmax, Consmax exhibits a reduced area footprint along with an increased maximum operating frequency. This finding is consistent across both open-source and proprietary EDA tools.
Furthermore, as depicted in Figure 9, an optimal energy consumption per cycle has been identified for each of the designs. At an operating frequency of 666 MHz, Consmax is projected to attain an energy efficiency of 0.2 pJ. When benchmarked against Softermax, this represents a significant energy saving of 3.5, and a 7.5 energy saving in comparison to Softermax ans Softmax respectively.
Figure 10 provides a visual representation of the power, area, and speed normalized for the three designs, evaluated at their respective peak operating frequencies. It is evident that Consmax achieves a superior speed while incurring lower power and area costs, underscoring its potential for highly efficient design applications.
@@ VI Conclusion
This paper presents ConSmax, a software-hardware co-design for efficient Softmax acceleration. ConSmax improves Softermax computing parallelism by avoiding the data synchronization, which produces maximum score and score summation. In addition, the bitwidth-split ConSmax hardware is lossless and scalable to calculating non-linear functions. The experiments show that ConSmax achieves a minuscule power consumption of 0.2 mW and area of 0.0008 mm at 1250-MHz working frequency and 16-nm CMOS technology. Compared to state-of-the-art Softmax hardware, ConSmax results in 3.35 power and 2.75 area savings with a comparable accuracy on GPT-2 model and the WikiText103 dataset.